A square loop of side 5cm being moved towards rite at a constant speed of 1cm/s d front edge enters d 20cm wide magnetic field at t=0 find d emf induced in d loop at a)t=2s b)t=10s c) t=22s  &  d) t=30s
here B = 0.6 T
 
A coil of radius 10cm & resistance 40ohm has 1000 turns it is placed with its plane vertical & its axis parallel to d magnetic meridian d coil is connected to a galvanometer & is rotated about d vertical diameter thru an angle of 180 find d charge which flows thru d galvanometer if d horizontal component of d earths magnetic field is BH = 3 x 10^-5 T
 
A rectangular frame of wire abcd has dimensions 32 x 8 cm & a total resistance of 2 ohm it is pulled out of a magnetic field B =0.02 T by applying a force of
3.2 x 10^-5 N it is found dat d frame moves with a constant speed find a) dis constant speed b)d emf induced in d loop , c) d potential difference between d points a &b & d) d potential difference btw d points c&d
 
d abov r H C VERMA problems & even some more quesn r der dey r
 
pg no 312 -- 80 , 81 , 82 , 85, 86
 
plz solve dis rates assured for d simplest method [even for a long method]
 


Grade:upto college level

1 Answers

ROSHAN MUJEEB
3 years ago
u = 1 cm/', B = 0.6T (a) At t = 2sec, distance moved = 2 × 1cm/s = 2cm (b) At t = 10sec distance moved = 10 × 1 = 10cm The flux linked does not change with time ∴ E = 0 (c) At t = 22sec distance = 22 × 1 = 22cm The loop is moving out of the field and 2cm outside. (d) At t = 30sec The loop is total outside and flux linked = 0 ∴ E = 0.